# Deep Learning-based Prediction of Clinical Trial Enrollment with Uncertainty Estimates

### v1

@@ 1 Introduction
Drug development is a complex process with multiple stages, including drug discovery, pre-clinical research, clinical research, review, and post-market safety monitoring. Each of these stages may include multiple steps. For instance, the clinical research evaluates potential treatments in different phases, which are referred to as clinical trials. Depending on the disease that is being addressed, a clinical trial can span months or even years, involving human participants. Additionally, clinical trials may be conducted in various sites distributed in multiple countries. Given the complexity of clinical trials, substantial investments and efforts are required for their successful execution, rendering the necessity of careful design of the trials. Even with meticulous planning, many clinical trials have failed to produce anticipated clinical outcomes. Recent analyses on different clinical indications have revealed that the success rate of clinical trials is approximately 7.97.97.9%, which is a high risk for pharmaceutical companies kim2023factors . One of the primary reasons for the failure of clinical trials is the insufficient enrollment of patients. Research shows that 191919% of the trials are terminated due to insufficient enrollment carlisle2015unsuccessful and 808080% of the trials do not meet initial enrollment goals, resulting in significant financial loss of up to \$888 million in revenue per day johnson2015evidence . Therefore, it is of utmost importance to be able to predict the clinical trial success in terms of enrollment before the trials start. This work focuses on enrollment prediction at planning phase.
@@ 2.2 Large Language Models
Large language models (LLMs) refer to state-of-the-art models with the capabilities of dealing with complex language understanding tasks such as question answering, machine translation, summarization, and text generation. LLMs typically possess an extensive number of parameters and necessitate huge computational resources for training and inference. The popularity of LLMs can be dated back to the introduction of GPT-3 brown2020language in 202220222022 when end-users were able to interact with LLMs via a chat interface. Since then, many LLMs including LLama dubey2024llama , Gemini team2024gemini , and Claude kevian2024capabilities have been introduced with emergent capabilities of solving diverse reasoning tasks and processing multi-modality data.
@@ 3.2 Deterministic Modeling
As discussed in Section 1, a clinical trial is generally complex, and has multiple attributes of miscellaneous types (see Table 1). In order to handle these attributes, similar to wang2022trial2vec , we first group the set of trial attributes into two subsets, namely Key and Context. The Key set contains important attributes that allow the positioning of the trial of interest among reference clinical trials. These attributes are: phase, country, therapeutic area (TA), sponsor, planned number of participants, and planned number of sites. We then apply pre-processing operations to these attributes using MultiLabelBinarizer [^1] https://scikit-learn.org/stable/modules/generated/sklearn.preprocessing.MultiLabelBinarizer.html and standard transformation (a.k.a. Z-score) to obtain two embeddings $\mathbf{x}^{cat}\in\mathbb{R}^{d_{1}}$ and $\mathbf{x}^{num}\in\mathbb{R}^{d_{2}}$. Conversely, the Context set is a mixture of textual and categorical attributes such as title, objective, mechanism of action, indication (disease), inclusion, and exclusion criteria. To effectively process this set, we first apply a simple text serialization technique fang2024large to form a unified text from attributes, that is we concatenate these attributes using a separator. We then generate an embedding $\mathbf{x}^{emb}\in\mathbb{R}^{d_{3}}$ for the unified text using Clinical Longformer li2022clinical . We select Clinical Longformer because it has the context length of 409640964096 tokens, which eliminates the issue of context being truncated given that the concatenated text may have thousands of tokens. More importantly, this model has been fine-tuned on clinical corpora, which is capable of capturing nuances in clinical text.
[…]
@@
$\mathcal{L}(\Theta)=\frac{1}{n}\sum_{i=1}^{n}\left|\text{ln}\ y_{i}-\text{ln}\ \hat{y_{i}}\right|,$ (8)
@@ 3.4 Deep Poisson-Gamma
$\rho_{i,s,t}=\begin{cases}\sim Poisson(\mu_{i,s})&\text{if }t\geq\theta_{i,s}\\ 0&\text{otherwise}\end{cases}\$ (12)
@@
$\Gamma_{\mu,i}\equiv\Gamma(\alpha_{\mu,i},\lambda_{\mu,i})\$ (16)
$\Gamma_{\theta,i}\equiv\Gamma(\alpha_{\theta,i},\lambda_{\theta,i})\$ (17)
@@
For the second approach, we implement a filtering-and-fitting baseline inspired by Anisimov anisimov2008using and Zhong et al. zhong2024enrollment . The method identifies, for each trial $C_{i}$, historical trials with similar characteristics (based on the categorical features listed in Table 1) and fits Gamma distributions $\Gamma_{\mu,i}$ and $\Gamma_{\theta,i}$ to, respectively, their historical site-level enrollment rates and startup times using maximum likelihood estimation (MLE).
@@ 4.1 Dataset
We collect trial-level data from two primary sources, namely IQVIA Data Query System (DQS [^2] https://www.iqvia.com/solutions/technologies/orchestrated-clinical-trials/planning-suite/data-query-system) and Citeline [^3] https://www.citeline.com/en. DQS contains proprietary operational data about clinical trials, principal investigators, and facilities. The data in DQS are provided by pharmaceutical companies or collected from shared data collaborations such as Investigator Registry or the Investigator Databank. Thanks to its volume and high quality, DQS enables sponsors and contract research organizations (CROs) to find the best matched investigators and facilities, and hence effectively support clinical trial planning. Data from Citeline come from more than 404040k sources [^4] https://www.informa.com/globalassets/documents/investor-relations/2019/investorday/citeline.pdf of monitored information, and are curated by their scientists. Similar to DQS, Citeline data contain information about trials, investigators, and facilities. Citeline data lack the detailed site-level performance metrics that DQS offers, however, Citeline provides extra features and metrics at trial and drug levels. Hence, we collect the information about number of sites, trial duration, and number of enrolled patients from DQS while other structured and textual attributes are taken from Citeline. Another data source is ClinicalTrials.gov, which is a publicly available database established by National Institute of Health (NIH). It should be noted that a large portion of Citeline trial data can be found in this database.
Our dataset contains more than 11.411.411.4k clinical trials of different therapeutic areas such as Oncology, Inflammation, and Cardiovascular. Since our method is based on supervised learning, we only consider trials with known outcome. For this reason, the status of a trial in our dataset must be either Completed or Closed. The number of enrolled patients may vary greatly across trials depending on the nature of therapeutic areas, ranging between 111 and 11.711.711.7k. A comprehensive list of attributes along with their corresponding summary descriptions is presented in Table 1.
Table 1: The list of attributes and label in our dataset collected from Citeline and DQS. The target enrollment is the planned number of patients set by clinical trial managers. The number of enrolled patients (No. enrolled) is the actual number of patients when the corresponding trial completes, which is our prediction target. It should be noted that the target enrollment number in some cases is close to the actual enrolled number but in many cases there are wide discrepancies between those numbers. “ - ” stands for attributes that are not applicable to the corresponding columns and TA stands for Therapeutic Area. Categorical attributes may be single-valued or multi-valued; those marked with a dagger are multi-valued. Attribute Phase has eight different values because some trials are multi-phase (e.g., II/III).
Attribute Source Type Min Median Mean Max Cardinality
Title Citeline Text - - - - -
Objective Citeline Text - - - - -
TA Citeline Categorical - - - - 999
Indication Citeline Categorical - - - - 970970970
Mechanism Citeline Categorical - - - - 995995995
Sponsor Citeline Categorical - - - - 962962962
Sponsor Type Citeline Categorical - - - - 101010
Country Citeline Categorical - - - - 147147147
Target enrollment Citeline Numeric 333 126126126 393.03393.03393.03 849658496584965 -
Drug name Citeline Categorical - - - - 341834183418
Inclusion Citeline Text - - - - -
Exclusion Citeline Text - - - - -
Phase DQS Categorical - - - - 8
No. site DQS Numeric 111 151515 42.5642.5642.56 114011401140 -
No. enrolled DQS Numeric 111 858585 292.24292.24292.24 117191171911719 -
@@ 4.2.1 Patient Enrollment Prediction Setting
In line with standard procedures, we split the data into three subsets, the training set (941094109410 trials), the development set (100010001000 trials), and the test set(100010001000 trials). The dataset is split in such a way that the distributions of number of enrolled patients in these training, development and test sets are similar. We then optimize the model parameters using the training set and select the model exhibiting the best performance on the development set. Subsequently, the evaluation metrics are calculated on the test set to obtain an unbiased assessment of the proposed model. We utilize two widely used metrics for measuring regression performance: the Mean Absolute Error (MAE) and the Coefficient of Determination ($R^{2}$). $R^{2}$ is popular metric for regression, however, its reliability can be compromised in case the data contain outliers. Hence, it is beneficial to complement $R^{2}$ with MAE to quantify the average magnitude of errors. It is important to note that although the model is trained on the logarithmic scale (see Section 3), MAE and $R^{2}$ are calculated on the original scale (i.e., number of participants) of data to maintain interpretability.
@@
We employ both classical machine learning models and advanced language models (PLMs/LLMs) as baselines for performance comparison. For classical models, we use the established gradient boosting models, including XGBoost chen2016xgboost and LightGBM ke2017lightgbm . These models have been known for very high performance on tabular datasets. In order to use these models, encoded attributes are concatenated to create a unified input embedding. For PLMs, we fine-tune BioBERT lee2020biobert , ClinicalBERT wang2023optimized , and Clinical Longformer li2022clinical for regression task; these models take as input the serialized text as discussed in Section 3.2. For LLMs, we employ the fine-tuning approach, leveraging Llama2 touvron2023llama . Similar to Longformer, Llama2 features an extensive context length of 409640964096 tokens, which are sufficiently ample to process long clinical text. Given the limited computational resources, we fine-tune the smallest version of Llama2 (7b) with the following instruction:
[…]
@@
Analyze the subject enrollment of the following clinical trial enclosed in square brackets. Determine the number of subjects as follows:
@@
{Clinical Trial Text}: {Number of subjects}. Using this format, we prepare the training dataset for LLama2 and fine-tune the model using LoRA technique hu2022lora . The implementation is based on library PEFT [^5] https://github.com/huggingface/peft.
@@
For training the proposed model, we leverage AdamW optimization algorithm loshchilov2017decoupled . We use a small dropout rate of 0.30.30.3 for the category branch but we do not apply dropout regularization on other branches. The batch size is set to 256256256. We also use two different learning rates: the same learning rate of 0.00010.00010.0001 for the input modules, while the learning rate for the rest of the model is 0.0010.0010.001.
@@ 4.2.2 Deep Poisson-Gamma Setting
To maintain the statistical stability and avoid irregularities of enrollment data, we choose to filter the data similar to the setting in Zhong et al. zhong2024enrollment . Specifically, we keep only trials with more than 101010 sites and a duration between 666 and 363636 months. We then use a subset of the training, development, and test sets described in Section 4.1 by filtering the trials according to the criteria described in Table 2. After filtering, the training, development, and test sets contain 164116411641, 152152152, and 174174174 trials, respectively.
Table 2: Trial selection criteria for the Poisson-Gamma modeling.
Attribute Value
Trial completion date Between 2015 and 2025
Number of sites More than 101010 sites
Trial duration Between 6 and 36 months
Trial phase II and III
We follow the same procedure outlined in Section 4.2.1, and introduce two additional metrics, namely, the Median Average Error (MedAE) and the 6-month coverage. The latter represents the percentage of trials for which the actual duration falls within a 6-month window centered around our prediction. For the first approach leveraging our stochastic model described in Section 3.4, we train our stochastic model with a learning rate of 0.00010.00010.0001 and a batch-size of 323232 for 256256256 epochs. After training our model $\mathcal{F}_{(\mu,\theta)}$, we use it to simulate the enrollment 102410241024 times with the Poisson-Gamma model. Each of these simulations produces a predicted duration $\hat{\tau}$. Subsequently, the predicted durations are averaged to obtain our final prediction.
For the “filtering-and-fitting” approach, we use an RMSprop optimizer to find the parameters $\alpha$ and $\lambda$ of our distributions $\Gamma_{\mu,i}$ and $\Gamma_{\theta,i}$ for each trial. The parameters of the optimizer are a learning rate of 0.010.010.01, a batch-size of 128128128 and the optimization is carried out across 512512512 epochs. The final prediction is obtained in the same way as for the first approach. Moreover, it should be noted that, for the two different approaches, the maximum duration allowed for the prediction is two times the trial duration upper-limit described in Table 2 to avoid unnecessary computation while keeping enough margin for potential errors.
@@ 4.3.1 Patient Enrollment Prediction
Table 3: The performance of the proposed method in comparison with baseline models. $R^{2}$ is the coefficient of determination. MAE is the mean absolute error. ($\uparrow$) indicates that higher values reflect better performance while ($\downarrow$) indicates that smaller values correspond to better performance. For XGBoost and LightGBM, * indicates that the embeddings of textual attributes are used. Baseline (first row) shows the MAE and $R^{2}$ calculated between planned and actual numbers.
Method $R^{2}$ ($\uparrow$) MAE ($\downarrow$)
Baseline (Target number of patients) 0.39 71.42
XGBoost* chen2016xgboost 0.70 66.74
XGBoost chen2016xgboost 0.72 62.22
LightGBM* ke2017lightgbm 0.71 62.61
LightGBM ke2017lightgbm 0.76 56.84
Llama2 touvron2023llama 0.070.070.07 110.98110.98110.98
ClinicalBERT* wang2023optimized 0.720.720.72 81.0881.0881.08
BioBERT* lee2020biobert $\mathbf{0.77}$ 69.3669.3669.36
Clinical-Longformer-153615361536li2022clinical 0.740.740.74 59.8859.8859.88
Clinical-Longformer-204820482048li2022clinical 0.730.730.73 59.6459.6459.64
Our Deterministic Model* 0.760.760.76 $\mathbf{51.99}$
Our Probabilistic Model* $\mathbf{0.77}$ 55.9255.9255.92
The performance of the proposed method compared to baseline models is presented in Table 3. Classical ML models perform reasonably good, surpassing the naive baseline of target number of patients, which achieves an $R^{2}$ of 0.390.390.39 and MAE of 71.4271.4271.42. It is also worth noting that simply concatenating textual embeddings to attribute embeddings does not generally improve performance. Notably, LightGBM without textual embeddings achieves top performance among classical ML models with an $R^{2}$ of 0.760.760.76 and MAE of 56.8456.8456.84. The PLM fine-tuning approach consistently achieves good $R^{2}$ scores. Nonetheless, ClinicalBERT and BioBERT do not have good performance on MAE, which is not the case for Clinical Longformer with an MAE of 59.6459.6459.64. This can be explained by the fact that Longformer has much longer context window compared to ClinicalBERT and BioBERT, which allows it to capture the details of the input clinical text.
Compared to the baselines, our models demonstrate superior performance in both metrics. Specifically, the deterministic model performs impressively, achieving an $R^{2}$ score of 0.760.760.76, which places it in the second-highest position among the models evaluated. The model excels in terms of MAE, delivering the best performance with a value of 51.9951.$9$ an improvement of approximately 999% over the leading baseline model. The stochastic model, despite its inherent variability, secures the second-best position in MAE and outperforms other models in $R^{2}$, achieving a score of 0.770.770.77. The superior performance of the proposed deterministic and stochastic models can be attributed to its architecture, which can effectively capture and represent multi-modality information.
@@
[…]

### v2

@@ 1 Introduction
Drug development is a complex process with multiple stages, including drug discovery, pre-clinical research, clinical research, review, and post-market safety monitoring. Each of these stages may include multiple steps. For instance, the clinical research evaluates potential treatments in different phases, which are referred to as clinical trials. Depending on the disease that is being addressed, a clinical trial can span months or even years, involving human participants. Additionally, clinical trials may be conducted in various sites distributed in multiple countries. Given the complexity of clinical trials, substantial investments and efforts are required for their successful execution, rendering the necessity of careful design of the trials. Even with meticulous planning, many clinical trials have failed to produce anticipated clinical outcomes. Recent analyses on different clinical indications have revealed that the success rate of clinical trials is approximately 7.97.9%, which is a high risk for pharmaceutical companies kim2023factors . One of the primary reasons for the failure of clinical trials is the insufficient enrollment of patients. Research shows that 1919% of the trials are terminated due to insufficient enrollment carlisle2015unsuccessful and 8080% of the trials do not meet initial enrollment goals, resulting in significant financial loss of up to \$88 million in revenue per day johnson2015evidence . Therefore, it is of utmost importance to be able to predict the clinical trial success in terms of enrollment before the trials start. This work focuses on enrollment prediction at planning phase.
@@ 2.2 Large Language Models
Large language models (LLMs) refer to state-of-the-art models with the capabilities of dealing with complex language understanding tasks such as question answering, machine translation, summarization, and text generation. LLMs typically possess an extensive number of parameters and necessitate huge computational resources for training and inference. The popularity of LLMs can be dated back to the introduction of GPT-3 brown2020language in $2022$ when end-users were able to interact with LLMs via a chat interface. Since then, many LLMs including LLama dubey2024llama , Gemini team2024gemini , and Claude kevian2024capabilities have been introduced with emergent capabilities of solving diverse reasoning tasks and processing multi-modality data.
@@ 3.2 Deterministic Modeling
As discussed in Section 1, a clinical trial is generally complex, and has multiple attributes of miscellaneous types (see Table LABEL:table:data_attribute). In order to handle these attributes, similar to wang2022trial2vec , we first group the set of trial attributes into two subsets, namely Key and Context. The Key set contains important attributes that allow the positioning of the trial of interest among reference clinical trials. These attributes are: phase, country, therapeutic area (TA), sponsor, planned number of participants, and planned number of sites. We then apply pre-processing operations to these attributes using MultiLabelBinarizer [^1] https://scikit-learn.org/stable/modules/generated/sklearn.preprocessing.MultiLabelBinarizer.html and standard transformation (a.k.a. Z-score) to obtain two embeddings $\mathbf{x}^{cat}\in\mathbb{R}^{d_{1}}$ and $\mathbf{x}^{num}\in\mathbb{R}^{d_{2}}$. Conversely, the Context set is a mixture of textual and categorical attributes such as title, objective, mechanism of action, indication (disease), inclusion, and exclusion criteria. To effectively process this set, we first apply a simple text serialization technique fang2024large to form a unified text from attributes, that is we concatenate these attributes using a separator. We then generate an embedding $\mathbf{x}^{emb}\in\mathbb{R}^{d_{3}}$ for the unified text using Clinical Longformer li2022clinical . We select Clinical Longformer because it has the context length of $4096$ tokens, which eliminates the issue of context being truncated given that the concatenated text may have thousands of tokens. More importantly, this model has been fine-tuned on clinical corpora, which is capable of capturing nuances in clinical text.
@@
[…]
@@ 3.4 Deep Poisson-Gamma
$\rho_{i,s,t}=\begin{cases}\sim Poisson(\mu_{i,s})&\text{if }t\geq\theta_{i,s}\\ 0&\text{otherwise}\end{cases}\\$ (12)
@@
$\Gamma_{\mu,i}\equiv\Gamma(\alpha_{\mu,i},\lambda_{\mu,i})\\$ (16)
$\Gamma_{\theta,i}\equiv\Gamma(\alpha_{\theta,i},\lambda_{\theta,i})\\$ (17)
@@
For the second approach, we implement a filtering-and-fitting baseline inspired by Anisimov anisimov2008using and Zhong et al. zhong2024enrollment . The method identifies, for each trial $C_{i}$, historical trials with similar characteristics (based on the categorical features listed in Table LABEL:table:data_attribute) and fits Gamma distributions $\Gamma_{\mu,i}$ and $\Gamma_{\theta,i}$ to, respectively, their historical site-level enrollment rates and startup times using maximum likelihood estimation (MLE).
@@ 4.1 Dataset
We collect trial-level data from two primary sources, namely IQVIA Data Query System (DQS [^2] https://www.iqvia.com/solutions/technologies/orchestrated-clinical-trials/planning-suite/data-query-system) and Citeline [^3] https://www.citeline.com/en. DQS contains proprietary operational data about clinical trials, principal investigators, and facilities. The data in DQS are provided by pharmaceutical companies or collected from shared data collaborations such as Investigator Registry or the Investigator Databank. Thanks to its volume and high quality, DQS enables sponsors and contract research organizations (CROs) to find the best matched investigators and facilities, and hence effectively support clinical trial planning. Data from Citeline come from more than 4040k sources [^4] https://www.informa.com/globalassets/documents/investor-relations/2019/investorday/citeline.pdf of monitored information, and are curated by their scientists. Similar to DQS, Citeline data contain information about trials, investigators, and facilities. Citeline data lack the detailed site-level performance metrics that DQS offers, however, Citeline provides extra features and metrics at trial and drug levels. Hence, we collect the information about number of sites, trial duration, and number of enrolled patients from DQS while other structured and textual attributes are taken from Citeline. Another data source is ClinicalTrials.gov, which is a publicly available database established by National Institute of Health (NIH). It should be noted that a large portion of Citeline trial data can be found in this database.
Our dataset contains more than 11.411.4k clinical trials of different therapeutic areas such as Oncology, Inflammation, and Cardiovascular. Since our method is based on supervised learning, we only consider trials with known outcome. For this reason, the status of a trial in our dataset must be either Completed or Closed. The number of enrolled patients may vary greatly across trials depending on the nature of therapeutic areas, ranging between $1$ and 11.711.7k. A comprehensive list of attributes along with their corresponding summary descriptions is presented in Table LABEL:table:data_attribute.
@@ 4.2.1 Patient Enrollment Prediction Setting
In line with standard procedures, we split the data into three subsets, the training set ($9410$ trials), the development set ($1000$ trials), and the test set($1000$ trials). The dataset is split in such a way that the distributions of number of enrolled patients in these training, development and test sets are similar. We then optimize the model parameters using the training set and select the model exhibiting the best performance on the development set. Subsequently, the evaluation metrics are calculated on the test set to obtain an unbiased assessment of the proposed model. We utilize two widely used metrics for measuring regression performance: the Mean Absolute Error (MAE) and the Coefficient of Determination ($R^{2}$). $R^{2}$ is popular metric for regression, however, its reliability can be compromised in case the data contain outliers. Hence, it is beneficial to complement $R^{2}$ with MAE to quantify the average magnitude of errors. It is important to note that although the model is trained on the logarithmic scale (see Section 3), MAE and $R^{2}$ are calculated on the original scale (i.e., number of participants) of data to maintain interpretability.
@@
We employ both classical machine learning models and advanced language models (PLMs/LLMs) as baselines for performance comparison. For classical models, we use the established gradient boosting models, including XGBoost chen2016xgboost and LightGBM ke2017lightgbm . These models have been known for very high performance on tabular datasets. In order to use these models, encoded attributes are concatenated to create a unified input embedding. For PLMs, we fine-tune BioBERT lee2020biobert , ClinicalBERT wang2023optimized , and Clinical Longformer li2022clinical for regression task; these models take as input the serialized text as discussed in Section 3.2. For LLMs, we employ the fine-tuning approach, leveraging Llama2 touvron2023llama . Similar to Longformer, Llama2 features an extensive context length of $4096$ tokens, which are sufficiently ample to process long clinical text. Given the limited computational resources, we fine-tune the smallest version of Llama2 (7b) with the following instruction:
@@
“ Analyze the subject enrollment of the following clinical trial enclosed in square brackets. Determine the number of subjects as follows:
@@
{Clinical Trial Text}: {Number of subjects}. ” Using this format, we prepare the training dataset for LLama2 and fine-tune the model using LoRA technique hu2022lora . The implementation is based on library PEFT [^5] https://github.com/huggingface/peft.
@@
For training the proposed model, we leverage AdamW optimization algorithm loshchilov2017decoupled . We use a small dropout rate of $0.3$ for the category branch but we do not apply dropout regularization on other branches. The batch size is set to $256$. We also use two different learning rates: the same learning rate of $0.0001$ for the input modules, while the learning rate for the rest of the model is $0.001$.
@@ 4.2.2 Deep Poisson-Gamma Setting
To maintain the statistical stability and avoid irregularities of enrollment data, we choose to filter the data similar to the setting in Zhong et al. zhong2024enrollment . Specifically, we keep only trials with more than $10$ sites and a duration between $6$ and $36$ months. We then use a subset of the training, development, and test sets described in Section 4.1 by filtering the trials according to the criteria described in Table LABEL:table:poisson_gamma_data_criteria. After filtering, the training, development, and test sets contain $635$, $53$, and $61$ trials, respectively.
We follow the same procedure outlined in Section 4.2.1, and introduce two additional metrics, namely, the Median Average Error (MedAE) and the 6-month coverage. The latter represents the percentage of trials for which the actual duration falls within a 6-month window centered around our prediction. For the first approach leveraging our stochastic model described in Section 3.4, we train our stochastic model with a learning rate of $0.0001$ and a batch-size of $128$ for $2048$ epochs. After training our model $\mathcal{F}_{(\mu,\theta)}$, we use it to simulate the enrollment $1024$ times with the Poisson-Gamma model. Each of these simulations produces a predicted duration $\hat{\tau}$. Subsequently, the predicted durations are averaged to obtain our final prediction.
For the “filtering-and-fitting” approach, we use an RMSprop optimizer to find the parameters $\alpha$ and $\lambda$ of our distributions $\Gamma_{\mu,i}$ and $\Gamma_{\theta,i}$ for each trial. The parameters of the optimizer are a learning rate of $0.01$, a batch-size of $128$ and the optimization is carried out across $512$ epochs. The final prediction is obtained in the same way as for the first approach. Moreover, it should be noted that, for the two different approaches, the maximum duration allowed for the prediction is two times the trial duration upper-limit described in Table LABEL:table:poisson_gamma_data_criteria to avoid unnecessary computation while keeping enough margin for potential errors.
@@ 4.3.1 Patient Enrollment Prediction
The performance of the proposed method compared to baseline models is presented in Table LABEL:table:node_classification_result. Classical ML models perform reasonably good, surpassing the naive baseline of target number of patients, which achieves an $R^{2}$ of $0.39$ and MAE of $71.42$. It is also worth noting that simply concatenating textual embeddings to attribute embeddings does not generally improve performance. Notably, LightGBM without textual embeddings achieves top performance among classical ML models with an $R^{2}$ of $0.76$ and MAE of $56.84$. The PLM fine-tuning approach consistently achieves good $R^{2}$ scores. Nonetheless, ClinicalBERT and BioBERT do not have good performance on MAE, which is not the case for Clinical Longformer with an MAE of $59.64$. This can be explained by the fact that Longformer has much longer context window compared to ClinicalBERT and BioBERT, which allows it to capture the details of the input clinical text.
Compared to the baselines, our models demonstrate superior performance in both metrics. Specifically, the deterministic model performs impressively, achieving an $R^{2}$ score of $0.76$, which places it in the second-highest position among the models evaluated. The model excels in terms of MAE, delivering the best performance with a value of $51.99$ - an improvement of approximately 99% over the leading baseline model. The stochastic model, despite its inherent variability, secures the second-best position in MAE and outperforms other models in $R^{2}$, achieving a score of $0.77$. The superior performance of the proposed deterministic and stochastic models can be attributed to its architecture, which can effectively capture and represent multi-modality information.
@@
The proposed stochastic model can be used for range estimation, in addition to point estimation. This ability enables the generation of confidence intervals around enrollment estimates, offering planners a range of plausible outcomes. Specifically, it is possible to predict an interval containing the number of patients with a confidence level of $(1-\alpha)100\%$, where $\alpha$ is the significance level. For instance, choosing $\alpha=0.1$ leads to a confidence level of $90\%$, a widely used confidence interval. Upon comparison with the actual number of patients, it is observed that $78.73\%$ of ground-truth values fall into the predicted $90\%$ confidence intervals, of which the median value of the interval widths is $99.35$. Adjusting the significance level directly influences the confidence level, interval width, and interval accuracy. This demonstrates the flexibility of the stochastic approach over the deterministic approach. Figure 4 illustrates the variation of confidence level and interval accuracy relative to interval width.